# Automating the Refinement of Reinforcement Learning Specifications

### v1

@@ 3.2 Correctness of AutoSpec
The following theorem establishes correctness of AutoSpec, showing that the specification $\phi_{r}$ computed by AutoSpec is indeed a refinement of the input specification $\phi$. The proof, provided in Appendix A.2, proceeds by proving that each of the four refinement procedures results in a specification refinement.
@@
AutoSpec provides soundness guarantees – as shown in Theorem 1, the produced specification is guaranteed to be a refinement of the original specification as in Definition 2, and every trajectory that satisfies the refined satisfaction must also satisfy the original specification. However, AutoSpec does not provide completeness guarantees. This is not a limitation of AutoSpec, but an inherent property of the specification refinement problem itself because the problem is undecidable. This is because, even for the simplest case of reachability specifications (e.g., “reach region $G$ with probability at least pp”), deciding whether a given policy satisfies a specification is undecidable for general continuous-state MDPs or probabilistic programs capable of encoding Turing-complete behavior kaminski2015hardness. Consequently, no specification refinement algorithm can be both sound and complete. AutoSpec therefore focuses on soundness, which is important towards ensuring that a policy for the refined task also solves the task defined by the original specification.
@@ 4.1 Experimental Setup
We integrate AutoSpec with two specification-guided RL algorithms: DiRL dirl, which uses Dijkstra-style graph search with systematic exploration, and LSTS lsts, which uses multi-armed bandits for edge selection with epsilon-greedy exploration. These algorithms differ fundamentally in their exploration strategies, allowing us to examine how AutoSpec’s effectiveness depends on the underlying learning algorithm.
We evaluate on two domains specifically chosen to stress-test different aspects of specification refinement:
n-Rooms: Grid-based navigation with walls and doors, providing controlled tests of specific failure modes. State space: $(x,y,\theta,d)\in\mathbb{R}^{4}$ (position, angle to goal, distance). Action space: $(v,\theta)\in\mathbb{R}^{2}$ (velocity, direction). The n-rooms domain has been extensively used in specification-guided RL research dirl; spectrl; vzikelic2024compositional as it provides clear geometric structure while still presenting challenging long-horizon tasks. Its modular room structure naturally creates the types of specification failures we aim to address: trap states at room boundaries, dangerous narrow passages between rooms, and multiple alternative paths through different door configurations.
PandaGym gallouedec2021pandagym: Robotic manipulation requiring 3D navigation around obstacles. This domain tests refinement in high-dimensional continuous control where geometric intuitions may not apply directly. Following recent work showing the challenges of specification-guided RL in manipulation tasks lsts, we use this domain to validate that our convex hull and hyperplane-based refinements remain effective in high-dimensional spaces where human intuition about specification failures is limited.
For learning edge policies, both algorithms use PPO DBLP:journals/corr/SchulmanWDRK17 with stable-baselines3 stable-baselines3 implementation, following the standard practice in recent specification-guided RL work dirl; vzikelic2024compositional. We use 2-layer networks (64 neurons each), learning rate 0.0003, and standard PPO hyperparameters. In all experiments we evaluate refinements using a deliberately high satisfaction threshold ($p=0.99$). The purpose of this choice is methodological: by selecting a probability level that is difficult to achieve under coarse or under-specified predicates, we can clearly observe how the cumulative probability of satisfying the specification improves as AutoSpec performs successive refinements. Using such a stringent threshold ensures that even small improvements in guidance become visible in the satisfaction curves and allows us to measure the full extent of the benefit provided by refinement, independent of how poorly the initial specification performs.
@@ 4.2 Algorithm-Dependent Effectiveness: DiRL vs LSTS
This specification structure, with multiple disjunctive branches and sequential compositions, represents the type of complex task decomposition that prior work spectrl; dirl has identified as necessary for real-world applications but challenging for existing algorithms. The 100-rooms scale specifically tests whether refinements remain effective when the state space is large enough that exhaustive exploration is infeasible, reflecting concerns raised in lsts about scalability of compositional methods.
@@
Evaluation on Randomized 100-rooms and predicate placement To evaluate the generalization capabilities of our framework, we deployed AutoSpec in a procedurally generated 4-Way Gridworld where wall connectivity, and predicate placement were fully randomized for each seed (see Appendix A.3.1 for generation details). This setup specifically tests the system’s ability to synthesize policies without reliance on hand-engineered specifications or environment-specific heuristics. As shown in Figure 6, AutoSpec significantly outperforms the DIRL baseline, achieving a terminal success probability of approximately $60\%$ compared to the baseline’s stagnation at $20\%$. These results confirm that AutoSpec autonomously identifies and resolves task bottlenecks, raising the success rate of critical transitions from $20\%$ to $90\%$ via automatic refinement. (Appendix  A.3.5))
@@ A.1 Refinement Algorithms
$e_{r}\leftarrow[u^{*}\rightarrow u,\textit{with }\beta(e_{r}):=\beta(e)]$ Redefine edge with new predicate
@@
  $e_{ui}\leftarrow[u\rightarrow u_{i},\textit{with }\beta(e_{r}):=\beta(e)]$ Define edges from source to parents
@@ A.2 Proof of Theorem 1
OrRefine. Suppose that $G_{r}=\textsc{OrRefine}(e,G,\zeta)$. OrRefine only adds edges between existing vertices in $G$. Specifically, for a problematic edge $e=u\rightarrow u^{\prime}$, it identifies existing edges $e_{i}=u_{i}\rightarrow u^{\prime}$ and adds new edges $e_{new}=u\rightarrow u_{i}$ where $\beta(e_{new})=\beta(e)$.
@@ A.3.3 Randomized Placement Strategy
    Start and Goal Initialization: We define the four corners of the grid as candidate zones. To ensure traversal complexity, the StartStart and GoalGoal regions are restricted to opposite corners. A pair of opposite corner zones (e.g., Top-Left and Bottom-Right) is selected uniformly at random. The StartStart region is assigned to one zone and the GoalGoal to the other, denoted as:
@@ A.3.6 Results for Individual Refinements
Figure 7: Evaluation of Reach Probabilities in the 9-Rooms Environment. (a) The layout of the 9-rooms environment, showing the walls, doors, and goal regions, and the estimated convex hull for the new reach region, showing how the refinement process effectively restructs the reachable states, leading to better satisfaction of the specification (b) A comparison of reach probabilities between DIRL dirl and the proposed AutoSpec approach. The x-axis denotes the number of steps, and the y-axis denotes the estimated probability of success.
@@
We conducted multiple experiments to validate our approach to refining coarse-grained SpectRL specifications for solving RL tasks. The goal of our experiments is to compare the performance of the original DiRL dirl algorithm and our integration of DiRL with AutoSpec, thus showcasing the ability of AutoSpec to refine SpectRL specifications that are challenging for the existing algorithms for RL from logical specifications. For learning edge policies, in both cases we use Proximal Policy Optimization (PPO) DBLP:journals/corr/SchulmanWDRK17, implemented using stable-baselines3 stable-baselines3. We employ a 2-layer neural network, each layer containing 64 neurons. We consider two environments.
@@
PandaGym. The Pandagym gallouedec2021pandagym reach environment has a robotic arm with an object picked up and the task is to place the object at the correct location. A wall blocking the path to the goal is invisible to the robot. The state space consists of the current position of the gripper arm in 3D and the goal position.

### v2

@@ 3.2 Correctness of AutoSpec
The following theorem establishes correctness of AutoSpec, showing that the specification $\phi_{r}$ computed by AutoSpec is indeed a refinement of the input specification $\phi$. The proof, provided in the Appendix, proceeds by proving that each of the four refinement procedures results in a specification refinement.
@@
AutoSpec provides soundness guarantees – as shown in Theorem 1, the produced specification is guaranteed to be a refinement of the original specification as in Definition 2, and every trajectory that satisfies the refined satisfaction must also satisfy the original specification. However, AutoSpec does not provide completeness guarantees. This is not a limitation of AutoSpec, but an inherent property of the specification refinement problem itself because the problem is undecidable. This is because, even for the simplest case of reachability specifications (e.g., “reach region $G$ with probability at least pp”), deciding whether a given policy satisfies a specification is undecidable for general continuous-state MDPs or probabilistic programs capable of encoding Turing-complete behavior Kaminski & Katoen (2015). Consequently, no specification refinement algorithm can be both sound and complete. AutoSpec therefore focuses on soundness, which is important towards ensuring that a policy for the refined task also solves the task defined by the original specification.
@@ 4.1 Experimental Setup
We integrate AutoSpec with two specification-guided RL algorithms: DiRL Jothimurugan et al. (2021), which uses Dijkstra-style graph search with systematic exploration, and LSTS Shukla et al. (2024), which uses multi-armed bandits for edge selection with epsilon-greedy exploration. These algorithms differ fundamentally in their exploration strategies, allowing us to examine how AutoSpec’s effectiveness depends on the underlying learning algorithm. We evaluate on two domains specifically chosen to stress-test different aspects of specification refinement:
n-Rooms: Grid-based navigation with walls and doors, providing controlled tests of specific failure modes. State space: $(x,y,\theta,d)\in\mathbb{R}^{4}$ (position, angle to goal, distance). Action space: $(v,\theta)\in\mathbb{R}^{2}$ (velocity, direction). The n-rooms domain has been extensively used in specification-guided RL research Jothimurugan et al. (2021; 2019); Zikelic et al. (2023b) as it provides clear geometric structure while still presenting challenging long-horizon tasks. Its modular room structure naturally creates the types of specification failures we aim to address: trap states at room boundaries, dangerous narrow passages between rooms, and multiple alternative paths through different door configurations.
PandaGym Gallouédec et al. (2021): Robotic manipulation requiring 3D navigation around obstacles. This domain tests refinement in high-dimensional continuous control where geometric intuitions may not apply directly. Following recent work showing the challenges of specification-guided RL in manipulation tasks Shukla et al. (2024), we use this domain to validate that our convex hull and hyperplane-based refinements remain effective in high-dimensional spaces where human intuition about specification failures is limited.
For learning edge policies, both algorithms use PPO Schulman et al. (2017) with stable-baselines3 Raffin et al. (2021) implementation, following the standard practice in recent specification-guided RL work Jothimurugan et al. (2021); Zikelic et al. (2023b). We use 2-layer networks (64 neurons each), learning rate 0.0003, and standard PPO hyperparameters. In all experiments we evaluate refinements using a deliberately high satisfaction threshold ($p=0.99$). The purpose of this choice is methodological: by selecting a probability level that is difficult to achieve under coarse or under-specified predicates, we can clearly observe how the cumulative probability of satisfying the specification improves as AutoSpec performs successive refinements. Using such a stringent threshold ensures that even small improvements in guidance become visible in the satisfaction curves and allows us to measure the full extent of the benefit provided by refinement, independent of how poorly the initial specification performs.
@@ 4.2 Algorithm-Dependent Effectiveness: DiRL vs LSTS
This specification structure, with multiple disjunctive branches and sequential compositions, represents the type of complex task decomposition that prior work Jothimurugan et al. (2019; 2021) has identified as necessary for real-world applications but challenging for existing algorithms. The 100-rooms scale specifically tests whether refinements remain effective when the state space is large enough that exhaustive exploration is infeasible, reflecting concerns raised in Shukla et al. (2024) about scalability of compositional methods.
@@
Evaluation on Randomized 100-rooms and predicate placement To evaluate the generalization capabilities of our framework, we deployed AutoSpec in a procedurally generated 4-Way Gridworld where wall connectivity, and predicate placement were fully randomized for each seed (see the Appendix for generation details). This setup specifically tests the system’s ability to synthesize policies without reliance on hand-engineered specifications or environment-specific heuristics. As shown in Figure 6, AutoSpec significantly outperforms the DIRL baseline, achieving a terminal success probability of approximately $60\%$ compared to the baseline’s stagnation at $20\%$. These results confirm that AutoSpec autonomously identifies and resolves task bottlenecks, raising success rate of critical transitions from $20\%$ to $90\%$ via automatic refinement, see the Appendix.
@@ A.1 Refinement Algorithms
$e_{r}\leftarrow[u^{*}\rightarrow u^{\prime},\textit{with }\beta(e_{r}):=\beta(e)]$ Redefine edge with new predicate
@@
  $e_{new}\leftarrow[u\rightarrow u_{i},\textit{with }\beta(e_{new}):=\beta(e)\text{ and }\beta(e_{i}):=\beta(e)\cap\beta(e_{i})]$ Define edges from source to parents
@@ A.2 Proof of Theorem 1
OrRefine. Suppose that $G_{r}=\textsc{OrRefine}(e,G,\zeta)$. OrRefine only adds edges between existing vertices in $G$. Specifically, for a problematic edge $e=u\rightarrow u^{\prime}$, it identifies existing edges $e_{i}=u_{i}\rightarrow u^{\prime}$ and adds new edges $e_{new}=u\rightarrow u_{i}$ where $\beta(e_{new})=\beta(e)$ and $\beta(e_{i})=\beta(e)\cap\beta(e_{i})$.
@@ A.3.3 Randomized Placement Strategy
    Start and Goal Initialization: We define the four corners of the grid as candidate zones. To ensure traversal complexity, the $Start$ and $Goal$ regions are restricted to opposite corners. A pair of opposite corner zones (e.g., Top-Left and Bottom-Right) is selected uniformly at random. The $Start$ region is assigned to one zone and the $Goal$ to the other, denoted as:
@@ A.3.6 Results for Individual Refinements
Figure 7: Evaluation of Reach Probabilities in the 9-Rooms Environment. (a) The layout of the 9-rooms environment, showing the walls, doors, and goal regions, and the estimated convex hull for the new reach region, showing how the refinement process effectively restructs the reachable states, leading to better satisfaction of the specification (b) A comparison of reach probabilities between DIRL Jothimurugan et al. (2021) and the proposed AutoSpec approach. The x-axis denotes the number of steps, and the y-axis denotes the estimated probability of success.
@@
We conducted multiple experiments to validate our approach to refining coarse-grained SpectRL specifications for solving RL tasks. The goal of our experiments is to compare the performance of the original DiRL Jothimurugan et al. (2021) algorithm and our integration of DiRL with AutoSpec, thus showcasing the ability of AutoSpec to refine SpectRL specifications that are challenging for the existing algorithms for RL from logical specifications. For learning edge policies, in both cases we use Proximal Policy Optimization (PPO) Schulman et al. (2017), implemented using stable-baselines3 Raffin et al. (2021). We employ a 2-layer neural network, each layer containing 64 neurons. We consider two environments.
@@
PandaGym. The Pandagym Gallouédec et al. (2021) reach environment has a robotic arm with an object picked up and the task is to place the object at the correct location. A wall blocking the path to the goal is invisible to the robot. The state space consists of the current position of the gripper arm in 3D and the goal position.